A small, spherical bead of mass 2.50 g is released from rest at t = 0 from a point under the surface of a viscous liquid. The terminal speed is observed to be vT = 1.98 cm/s.

(a) Find the value of the constant b in the equation R with arrow = −bv with arrow.

______________________________________N·s/m

(b) Find the time t at which the bead reaches 0.632vT.

__________________s

(c) Find the value of the resistive force when the bead reaches terminal speed.

__________________________N
Answer #1

(. t = 0.00202 57 t -os er 14* 2.02 ms Sol:- given m=2.54-0.0025 Terminal spoed (V1)= 1.98 emls Vy = 0.0178 m/s a) The forces